What is the equation in slope intercept form?
2x+7y-14=0

Answer by algebrahouse.com(1659)

"What is the equation in slope intercept form?
2x+7y-14=0"

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

2x + 7y - 14 = 0
-7y = 2x - 14 {subtracted 7y from both sides}
y = (-2/7)x + 2 {divided both sides by -7 to get in slope-intercept form}
www.algebrahouse.com